Select the correct choice that completes the sentence below. y = In xif and only if

answer options:

y=e^x

x=10^y

x=e^y

y=10^x
Answer #1

Please check out my solution.

y = ln (x) { here y equals logarithm of x to the base e }

=> y = log (x) { here base is e }

=> e^y = x { log (a) to the base b equals c then a=b^c }

Therefore the correct option is x=e^y (ANSWER)

HOPE IT HELPED :-)

please give it a thumbs up if you liked it.